OpenStudy (anonymous):

9x + y = 0 18x + 4y = –8

OpenStudy (anonymous):

What we trying to do dessie91?

OpenStudy (anonymous):

You can either use substitution or elimination here. Substitution: y=-9x 18x+4(-9x)=-8 -18x=-8 x=4/9 9x+y=0 9(4/9)+y=0 4+y=0 y=-4 For elimination: (9x+y=0) <---Multiply this by -2 to eliminate x and solve for y 18x+4y=-8 -18x-2y=0 18x+4y=-8 Add these equations together; you get: 2y=-8 y=-4 Plug -4 back into an equation and solve for x: 9x-4=0 9x=4 x=4/9